### v1

# Infinite families of irreducible polynomials over finite fields

Abstract.
Let $p$ be a prime number and $q$ a power of $p$. Let ${\mathbb{F}}_{q}$ be the finite field with $q$ elements. For a positive integer $n$ and a polynomial $\varphi(X)\in{\mathbb{F}}_{q}[X]$, let $d_{n,\varphi}(X)$ denote the denominator of the $n$th iterate of $\frac{1}{\varphi(X)}$. The polynomial $\varphi(X)$ is said to be inversely stable over ${\mathbb{F}}_{q}$ if all polynomials $d_{n,\varphi}(X)$ are irreducible polynomial over ${\mathbb{F}}_{q}$ and distinct. In this paper, we characterize a class of inversely stable polynomials over ${\mathbb{F}}_{q}$. Actually, for $\varphi(X)=X^{p}-X+\xi\in{\mathbb{F}}_{q}[X]$, we provide a sufficient and necessary condition for $\varphi(X)$ to be inversely stable over ${\mathbb{F}}_{q}$.
Key words and phrases:
Finite fields, Irreducible polynomials, Iterations of rational functions.
1991 Mathematics Subject Classification:
Primary 11T06
1. Introduction
Constructing irreducible polynomials over finite fields is a central problem in finite field theory. Over the past few decades, numerous approaches have been developed to address this issue, particularly focusing on compositions and iterations of polynomials. One approach explores finding polynomials $f(X)$ and $g(X)$ such that $f(g(X))$ is an irreducible polynomial (see [2, 5, 6, 7, 16]). Another direction of research is identifying polynomials $h$ such that all of its iterates remain irreducible over a given finite field. Such polynomials are called stable over the finite field. Despite considerable attention to this topic (see [1, 8, 14, 15]), results remain sparse and the problem is far from being completely understood. For further discussions on different constructions of irreducible polynomials, one can refer to [3, 9, 10, 11, 12, 13] .
Let $q$ be a power of a prime number $p$, and let ${\mathbb{F}}_{q}$ denote the finite field of order $q$. Consider a rational function $\Phi(x)=\frac{1}{\varphi(X)}$, where $\varphi(X)\in{\mathbb{F}}_{q}[X]$. Denote by $\Phi^{(n)}(X)$ the $n$th iterate of $\Phi(X)$, i.e.,
$\Phi^{(0)}(X)=X,\ \Phi^{(n)}(X)=\Phi(\Phi^{(n-1)}(X)),\ n=1,2,\ldots.$
Let $d_{n,\varphi}(X)$ be the denominator of $\Phi^{(n)}(X)$. We say that $\varphi(X)$ is an inversely stable polynomial over ${\mathbb{F}}_{q}$ if all polynomials in the sequence $\{d_{n,\varphi}\}_{n=1}^{\infty}$ are irreducible over ${\mathbb{F}}_{q}$ and are distinct.
Recently, the author [4] identified a class of inversely stable polynomials over the prime field ${\mathbb{F}}_{p}$, and proposed the following problem.
Problem 1.1.
Let $q=p^{e}$ with $e$ a positive integer. Characterize the conditions under which the polynomial $f(X)=X^{p^{t}}+aX+b\in{\mathbb{F}}_{q}[X]$, where $t$ is a positive integer, is inversely stable over ${\mathbb{F}}_{q}$.
In the case $e=t=1$, the author provided a complete solution to Problem 1.1 (see [4, Theorem 1.1]). However, for $e=t\geq 2$, it was shown that $f(X)=X^{p^{t}}+aX+b\in{\mathbb{F}}_{q}[X]$ is never inversely stable over ${\mathbb{F}}_{q}$. While a full solution to the general case remains elusive, this paper offers a partial solution for the case $e>t=1$.
Let ${\rm Tr}_{{\mathbb{F}}_{q}}$ be the absolute trace function from ${\mathbb{F}}_{q}$ to ${\mathbb{F}}_{p}$. For an element $\xi\in{\mathbb{F}}_{q}$ with ${\rm Tr}_{{\mathbb{F}}_{q}}(\xi)\neq 0$, define three sequences $\{a_{n}\},\{c_{n}\}$ and $\{d_{n}\}$ of ${\mathbb{F}}_{q}$ as follows:
$a_{1}=\xi,c_{1}=1,d_{1}=0;a_{2}=-1,c_{2}=\xi,d_{2}=-1;$
and for $n\geq 2$,
$\begin{cases}a_{n+1}=-a_{n}d_{n},\\ c_{n+1}=c_{n}^{2}(\xi-(c_{n}^{-1}d_{n})^{p}+c_{n}^{-1}d_{n}),\\ d_{n+1}=-c_{n}^{2}.\end{cases}$ (1.1)
Since none of the elements in $\{c_{n}\}$ is zero (as ${\rm Tr}_{{\mathbb{F}}_{q}}(\xi)\neq 0$), the sequences are well defined. Now we present the main result of this paper.
Theorem 1.2.
Let $p$ be a prime, and $q=p^{e}$ with $e$ a positive integer. Let $g(X)=X^{p}-X+\xi\in{\mathbb{F}}_{q}[X]$ with ${\rm Tr}_{{\mathbb{F}}_{q}}(\xi)\neq 0$. Then, $g(X)$ is inversely stable over ${\mathbb{F}}_{q}$ if and only if ${\rm Tr}_{{\mathbb{F}}_{q}}(a_{n}c_{n}^{-1})\neq 0$ for every positive integer $n$, where the sequences $\{a_{n}\}$ and $\{c_{n}\}$ are defined as in (1.1).
Clearly, for a given polynomial $g(X)=X^{p}-X+\xi\in{\mathbb{F}}_{q}[X]$ with ${\rm Tr}_{{\mathbb{F}}_{q}}(\xi)\neq 0$, Theorem 1.2 provides an efficient criterion to determine whether $g(X)$ is inversely stable over ${\mathbb{F}}_{q}$. The conditions presented in the theorem allow straightforward verification of inverse stability by examining the absolute traces of specific sequences derived from the polynomial $g(X)$. Here are some examples.
Example 1.3.
Let $p$ be a prime, and $q=p^{e}$ with $e$ a positive integer. Let $\xi$ be any nonzero element of ${\mathbb{F}}_{p}$. It is evident that all the terms in the sequences $\{a_{n}\}$, $\{c_{n}\}$ and $\{d_{n}\}$ lie in ${\mathbb{F}}_{p}$. We then know that $c_{1}=1$, and $c_{n+1}=\xi c_{n}^{2}$ for any $n\geq 2$, which leads to the expression $c_{n}=\xi^{2^{n-1}-1}$ for all $n\geq 2$. Consequently, $d_{n}=-c_{n-1}^{2}=-\xi^{2^{n-1}-2}$ for any $n\geq 2$. Furthermore, using the recurrence relation $a_{n+1}=-a_{n}d_{n}$, $n=2,3,\ldots$, we compute that
$a_{n}=\frac{a_{n}}{a_{n-1}}\cdot\frac{a_{n-1}}{a_{n-2}}\cdots\frac{a_{3}}{a_{2}}\cdot a_{2}=(-d_{n-1})\cdot(-d_{n-2})\cdots(-d_{2})\cdot a_{2}=-\xi^{2^{n-1}-2n+2}$
for any $n\geq 3$. Therefore, we derive that $a_{n}c_{n}^{-1}=\xi$ if $n=1$, and $a_{n}c_{n}^{-1}=-\xi^{3-2n}$ if $n\geq 2$. Thus, we have
${\rm Tr}_{{\mathbb{F}}_{q}}(a_{n}c_{n}^{-1})=\begin{cases}e\xi,&\text{if}\ n=1,\\ -e\xi^{3-2n},&\text{if}\ n\geq 2.\end{cases}$
Therefore, by Theorem 1.2, $g(X)$ is inversely stable over ${\mathbb{F}}_{q}$ if and only if $p\nmid e$.
Example 1.4.
Let $p=3$ and $q=9$. Take $\xi=w$, where $w$ is a primitive element of $\mathbb{F}_{9}$ and $X^{2}+2X+2$ is the minimal polynomial over $\mathbb{F}_{3}$. By direct computations, we have the following:
$a_{1}=w,c_{1}=1,d_{1}=0,a_{1}c_{1}^{-1}=w,{\rm Tr}_{\mathbb{F}_{9}}(a_{1}c_{1}^{-1})=1,$
$a_{2}=2,c_{2}=w,d_{2}=2,a_{2}c_{2}^{-1}=2w+1,{\rm Tr}_{\mathbb{F}_{9}}(a_{2}c_{2}^{-1})=1,$
and for $n\geq 1$
$\begin{cases}a_{3n}=2,c_{3n}=2w,d_{3n}=2w+2,a_{3n}c_{3n}^{-1}=w+2,{\rm Tr}_{\mathbb{F}_{9}}(a_{3n}c_{3n}^{-1})=2,\\ a_{3n+1}=2w+2,c_{3n+1}=2w+2,d_{3n+1}=2w+2,a_{3n}c_{3n}^{-1}=1,{\rm Tr}_{\mathbb{F}_{9}}(a_{3n+1}c_{3n+1}^{-1})=2,\\ a_{3n+2}=1,c_{3n+2}=2w,d_{3n+2}=1,a_{3n+2}c_{3n+2}^{-1}=2w+1,{\rm Tr}_{\mathbb{F}_{9}}(a_{3n+2}c_{3n+2}^{-1})=1.\end{cases}$
Thus, from Theorem 1.2 we conclude that $X^{3}-X+w$ is an inversely stable polynomial over $\mathbb{F}_{9}$.
Example 1.5.
Let $p=5$ and $q=25$. Take $\xi=w$, where $w$ is a primitive element of $\mathbb{F}_{25}$ and $X^{2}+4X+2$ is the minimal polynomial over $\mathbb{F}_{5}$. It is computed that $a_{8}=3w+3$, $c_{8}=w$, $a_{8}c_{8}^{-1}=w+2$ and ${\rm Tr}_{\mathbb{F}_{25}}(a_{8}c_{8}^{-1})=0$. Therefore, by Theorem 1.2, we conclude that $X^{5}-X+w$ is not inversely stable over $\mathbb{F}_{25}$.
The structure of the paper is as follows: In Section 2, we introduce a series of lemmas that lay the groundwork for the main results. Section 3 is devoted to proving Theorem 1.2.
2. Lemmas
In this section, we introduce several useful lemmas that will form the foundation for proving our main results. Let $K$ be a finite field, and $L$ be an extension field of degree $m$. For $\alpha\in L$, denote by ${\rm Tr}_{L/K}(\alpha)$ the trace of $\alpha$ from $L$ to $K$, that is,
${\rm Tr}_{L/K}(\alpha)=\alpha+\alpha^{|K|}+\alpha^{|K|^{2}}+\cdots+\alpha^{|K|^{m-1}},$
where $|K|$ is the order of $K$. Additionally, the absolute trace of $\alpha$, denoted as ${\rm Tr}_{L}(\alpha)$, is defined by ${\rm Tr}_{L/{\mathbb{F}}_{p}}(\alpha)$, where ${\mathbb{F}}_{p}$ is the prime field of $K$.
Throughout this section, we assume that $K$ is a finite field with characteristic $p$. The first two lemmas follow immediately from basic properties of traces and irreducible polynomials.
Lemma 2.1.
[17, Equation (2.2)] Let $\alpha$ be an element of an extension field of $K$. Then the coefficient of the second highest term of the minimal polynomial of $\alpha$ over $K$ is equal to $-{\rm Tr}_{K(\alpha)/K}(\alpha)$.
Lemma 2.2.
Let $f(X)$ be a polynomial over $K$ of degree $m$. Define $g(X):=X^{m}f(\frac{1}{X})$. Then $f(X)$ is irreducible over $K$ if and only if $g(X)$ is irreducible over $K$.
Lemma 2.3.
Let $f(X)=X^{p}-X+\xi\in K[X]$ be irreducible over $K$, and let $\gamma$ be a root in an extension field of $K$. For any $a,b,c,d\in K$ with $c$ and $d$ not both zero, the following hold:
  1. (a)
    If $c=0$, then ${\rm Tr}_{K(\gamma)/K}(\frac{a\gamma+b}{c\gamma+d})=0$ if $p\geq 3$, and ${\rm Tr}_{K(\gamma)/K}(\frac{a\gamma+b}{c\gamma+d})=\frac{a}{d}$ if $p=2$.
  2. (b)
    If $c\neq 0$, then
    ${\rm Tr}_{K(\gamma)/K}\left(\frac{a\gamma+b}{c\gamma+d}\right)=\frac{bc-ad}{c^{2}(\xi-(\frac{d}{c})^{p}+\frac{d}{c})}.$
Proof.
The result of Item (a) follows immediately from the linearity of trace and Lemma 2.1. For Item (b), assume $c\neq 0$. Using the linearity of trace, one deduces that
${\rm Tr}_{K(\gamma)/K}\left(\frac{a\gamma+b}{c\gamma+d}\right)$ $={\rm Tr}_{K(\gamma)/K}\left(\frac{a}{c}+\frac{bc^{-1}-adc^{-2}}{\gamma+c^{-1}d}\right)$
$={\rm Tr}_{K(\gamma)/K}\left(\frac{a}{c}\right)+{\rm Tr}_{K(\gamma)/K}\left(\frac{bc^{-1}-adc^{-2}}{\gamma+c^{-1}d}\right)$
$=c^{-2}(bc-ad){\rm Tr}_{K(\gamma)/K}\left(\frac{1}{\gamma+c^{-1}d}\right).$ (2.1)
Note that $f(X)$ is the minimal polynomial of $\gamma$ over $K$. It implies that
$f(X-c^{-1}d)=(X-c^{-1}d)^{p}-(X-c^{-1}d)+\xi=X^{p}-X+\xi-\left(\frac{d}{c}\right)^{p}+\frac{d}{c}$
is the minimal polynomial of $\gamma+c^{-1}d$ over $K$. By Lemma 2.2 we then know that
$X^{p}f\left(\frac{1}{X}-c^{-1}d\right)=\left(\xi-\left(\frac{d}{c}\right)^{p}+\frac{d}{c}\right)X^{p}-X^{p-1}+1$
is irreducible over $K$, and then $X^{p}-(\xi-(\frac{d}{c})^{p}+\frac{d}{c})^{-1}X^{p-1}+(\xi-(\frac{d}{c})^{p}+\frac{d}{c})^{-1}$ is the minimal polynomial of $(\gamma+c^{-1}d)^{-1}$ over $K$. It follows from Lemma 2.1 that
${\rm Tr}_{K(\gamma)/K}\left(\frac{1}{\gamma+c^{-1}d}\right)=\left(\xi-\left(\frac{d}{c}\right)^{p}+\frac{d}{c}\right)^{-1}.$
Putting this into (2), we derive the desired result. ∎
Lemma 2.4.
[17, Corollary 3.79] For $f(X)=X^{p}-X+\xi\in K[X]$, we have that $f(X)$ is irreducible over $K$ if and only if ${\rm Tr}_{K}(\xi)\neq 0$.
Lemma 2.5.
Let $\overline{K}$ be the algebraic closure of $K$. Let $\infty$ be the point at infinity, i.e., $\infty=\frac{1}{0}$. For a polynomial $g(X)=X^{p}-X+\xi\in K[X]$, consider the rational map $G(X):=\frac{1}{g(X)}$ from $\overline{K}\cup\{\infty\}$ to itself. Then both of the following hold.
  1. (a)
    The map $G(X)$ is $p$-to-one, except that the preimage of 00 is $\infty$.
  2. (b)
    If ${\rm Tr}_{K}(\xi)\neq 0$, then $d_{n,g}(X)$ is a polynomial over $K$ of degree $p^{n}$ for every positive integer $n$, where $d_{n,g}(X)$ is the denominator of the $n$th iterate $G^{(n)}(X)$ of $G(X)$
Proof.
First, let us prove Item (a). Clearly, for the map $G(X)$ the preimage of 00 is $\infty$. Now let $\gamma$ be any fixed element of {}$\{\infty\}\cup\overline{K}\setminus\{0\}$. Note that the polynomial $g(X)-\frac{1}{\gamma}$ has no multiple roots, and then $g(X)-\frac{1}{\gamma}$ has exactly $p$ distinct roots in $\overline{K}$. Thus, the preimage set of $\gamma$ has order $p$, and the result of Item (a) follows.
Next, we prove Item (b). Let $n$ be a given positive integer. It is clear that $\deg(d_{n,g}(X))\leq p^{n}$. Now we show that $\deg(d_{n,g}(X))\geq p^{n}$, completing the proof of Item (b). To do this, we define $n$ sets:
[…]
We claim that $\infty\not\in(G^{(i)})^{-1}(\infty)$ for any $1\leq i\leq n$. Suppose that $\infty\in(G^{(i)})^{-1}(\infty)$ for some $i$, and let $k$ be the smallest positive integer with this property, that is,
$G^{(1)}(\infty)\neq\infty,\ldots,G^{(k-1)}(\infty)\neq\infty,G^{(k)}(\infty)=\infty.$
Let $\alpha_{j}:=G^{(j)}(\infty)$ for each $j$ with $1\leq j\leq k$. Then $\alpha_{1},\ldots,\alpha_{k-1}$ are elements in $K$, and $g(\alpha_{k-1})=0$, implying that $g(X)$ has a root of $K$. However, by Lemma 2.4, $g(X)$ is irreducible over $K$ when ${\rm Tr}_{K}(\xi)\neq 0$, a contradiction. Therefore, $\infty\not\in(G^{(i)})^{-1}(\infty)$ for $i$. Now applying the result from Item (a), we see that
#$\#G^{(1)}(\infty)=p,\ \text{and}\ \#G^{(i+1)}(\infty)=p\#G^{(i)}(\infty)\ \text{for\ all}\ 1\leq i\leq n-1.$
Thus, #$\#G^{(n)}(\infty)=p^{n}$, which implies that $d_{n,g}(X)$ has exactly $p^{n}$ distinct roots in $\overline{K}$. Therefore, $\deg(d_{n,g}(X))\geq p^{n}$, completing the proof of Item (b).
This concludes the proof of Lemma 2.5. ∎
3. Proof of Theorem 1.2
In this section, we primarily provide the proof of Theorem 1.2.
Proof of Theorem 1.2. Let $g(X)=X^{p}-X+\xi\in{\mathbb{F}}_{q}[x]$ with ${\rm Tr}_{{\mathbb{F}}_{q}}(\xi)\neq 0$. Let $\{a_{n}\},\{c_{n}\}$ and $\{d_{n}\}$ be the sequences over ${\mathbb{F}}_{q}$ defined as in (1.1). Let $G(X)=\frac{1}{g(X)}$, and for any positive integer $n$, let $G^{(n)}(X)$ be the $n$th iterate of $G(X)$. The denominator of $G^{(n)}(X)$ is denoted by $d_{n,g}(X)$. We aim to prove that $g(X)$ is inversely stable over ${\mathbb{F}}_{q}$ if and only if ${\rm Tr}_{{\mathbb{F}}_{q}}(a_{n}c_{n}^{-1})\neq 0$ for every positive integer $n$.
First, we prove the necessity. Let $g(X)=X^{p}-X+\xi$ be an inversely stable polynomial over ${\mathbb{F}}_{q}$. We are going to prove that ${\rm Tr}_{{\mathbb{F}}_{q}}(a_{n}c_{n}^{-1})\neq 0$ for every positive integer $n$. Let $n$ be any fixed positive integer. Note that $d_{n,g}(X)$ is irreducible over ${\mathbb{F}}_{q}$, and it is of degree $p^{n}$ from Lemma 2.5. Let $\beta_{n}$ be a root of $d_{n,g}(X)$ in the algebraic closure of ${\mathbb{F}}_{q}$. Then ${\mathbb{F}}_{q}(\beta_{n})$ is an extension field of degree $p^{n}$. For every integer $i$ with $11\leq i\leq n-$ 1, define $\beta_{i}:=G^{(n-i)}(\beta_{n})$. One then readily finds that
$\beta_{1}^{p}-\beta_{1}+\xi=0,\ \text{and}$ (3.1)
$\beta_{i}^{p}-\beta_{i}+\frac{\xi\beta_{i-1}-1}{\beta_{i-1}}=0\ \text{for\ any}\ 2\leq i\leq n.$ (3.2)
Now define a sequence $\{K_{i}\}_{i=0}^{n}$ of finite fields by
$K_{0}={\mathbb{F}}_{q},\ \text{and}\ K_{i}=K_{i-1}(\beta_{i})\ \text{for}\ 1\leq i\leq n.$
Let $h_{1}(X)=X^{p}-X+\xi$, and for $2\leq j\leq n$, $h_{j}(X)=X^{p}-X+\frac{\xi\beta_{j-1}-1}{\beta_{j-1}}$. Clearly, $h_{i}(X)\in K_{i-1}[X]$ and $h_{i}(\beta_{i})=0$ for any $1\leq j\leq n$. It implies that $[K_{i}:K_{i-1}]\leq p$ for each $1\leq i\leq n$. So, we obtain that
$p^{n}=[{\mathbb{F}}_{q}(\beta_{n}):{\mathbb{F}}_{q}]=[K_{n}:K_{0}]=\prod_{i=1}^{n}[K_{i}:K_{i-1}]\leq p^{n},$
and then $[K_{i}:K_{i-1}]=p$ for any $1\leq i\leq n$. It follow that $h_{i}(X)$ is irreducible over $K_{i-1}$ for each $1\leq i\leq n$. In particular, $h_{n}(X)=X^{p}-X+\frac{\xi\beta_{n-1}-1}{\beta_{n-1}}$ is irreducible over $K_{n-1}$. Applying Lemma 2.4, we then have
${\rm Tr}_{K_{n-1}}\left(\frac{\xi\beta_{n-1}-1}{\beta_{n-1}}\right)\neq 0.$
On the other hand, noting that each $h_{i}(X)$ is irreducible over $K_{i-1}$, and employing Lemma 2.3 again and again, we compute that
${\rm Tr}_{{K_{n-1}}/K_{n-2}}\left(\frac{\xi\beta_{n-1}-1}{\beta_{n-1}}\right)=\frac{a_{2}\beta_{n-2}}{c_{2}\beta_{n-2}+d_{2}},$
${\rm Tr}_{{K_{n-2}}/K_{n-3}}\left(\frac{a_{2}\beta_{n-2}}{c_{2}\beta_{n-2}+d_{2}}\right)=\frac{a_{3}\beta_{n-3}}{c_{3}\beta_{n-3}+d_{3}},$
$\vdots$
${\rm Tr}_{{K_{2}}/K_{1}}\left(\frac{a_{n-2}\beta_{2}}{c_{n-2}\beta_{2}+d_{n-2}}\right)=\frac{a_{n-1}\beta_{1}}{c_{n-1}\beta_{1}+d_{n-1}},$
${\rm Tr}_{{K_{1}}/K_{0}}\left(\frac{a_{n-1}\beta_{1}}{c_{n-1}\beta_{1}+d_{n-1}}\right)=a_{n}c_{n}^{-1}.$
Therefore, by the transitivity of trace we arrive at
${\rm Tr}_{K_{0}}(a_{n}c_{n}^{-1})={\rm Tr}_{K_{n-1}}\left(\frac{\xi\beta_{n-1}-1}{\beta_{n-1}}\right)\neq 0.$
The necessity is proved.
Next, we focus on the sufficiency. Let $\xi\in{\mathbb{F}}_{q}$ satisfy that ${\rm Tr}_{{\mathbb{F}}_{q}}(a_{n}c_{n}^{-1})\neq 0$ for any $n\geq 1$. We shall prove that $g(X)=X^{p}-X+\xi$ is inversely stable over ${\mathbb{F}}_{q}$. Let $n$ be any given positive integer. By Lemma 2.5, we know that $d_{n,g}(X)$ is of degree $p^{n}$, so we only need to prove that $d_{n,g}(X)$ is irreducible over ${\mathbb{F}}_{q}$, which will be done in what follows. Let $\beta_{i}$, $K_{i}$ and $h_{i}(X)$ be defined as in the proof of necessity. First, we show that $h_{i}(X)$ is irreducible over $K_{i-1}$ for any $1\leq i\leq n$. Obviously, $h_{1}(X)$ is irreducible over $K_{0}$, and we need to prove that $h_{i}(X)$ is irreducible over $K_{i-1}$ for any $2\leq i\leq n$, which is equivalent to that
${\rm Tr}_{K_{i-1}}\left(\frac{\xi\beta_{i-1}-1}{\beta_{i-1}}\right)\neq 0$ (3.3)
for any $2\leq i\leq n$. Now we prove (3.3) by mathematical induction on $i$.
  • For $i=2$, noting that $h_{1}(X)$ is irreducible over $K_{0}$, $h_{1}(\beta_{1})=0$ and $K_{1}=K_{0}(\beta)$, by using Lemma 2.3 one then checks that
    ${\rm Tr}_{K_{1}/K_{0}}\left(\frac{\xi\beta_{1}-1}{\beta_{1}}\right)=a_{2}c_{2}^{-1},$
    and then
    ${\rm Tr}_{K_{1}}\left(\frac{\xi\beta_{1}-1}{\beta_{1}}\right)={\rm Tr}_{K_{0}}(a_{2}c_{2}^{-1})\neq 0.$
    It is to say that (3.3) is true for $i=2$.
  • Assume that (3.3) is true for all $i$ with $i\leq s-1$. It implies from Lemma 2.4 that $h_{i}(X)$ is irreducible over $K_{i-1}$ for each $11\leq i\leq s-$ 1. Since $K_{i}=K_{i-1}(\beta_{i})$ and $h_{i}(\beta_{i})=0$ for any $11\leq i\leq s-$ 1, by using Lemma 2.3 repeatedly we derive that
    ${\rm Tr}_{{K_{s-1}}/K_{s-2}}\left(\frac{\xi\beta_{s-1}-1}{\beta_{s-1}}\right)=\frac{a_{2}\beta_{s-2}}{c_{2}\beta_{s-2}+d_{2}},$
    ${\rm Tr}_{{K_{s-2}}/K_{s-3}}\left(\frac{a_{2}\beta_{s-2}}{c_{2}\beta_{s-2}+d_{2}}\right)=\frac{a_{3}\beta_{s-3}}{c_{3}\beta_{s-3}+d_{3}},$
    $\vdots$
    ${\rm Tr}_{{K_{2}}/K_{1}}\left(\frac{a_{s-2}\beta_{2}}{c_{s-2}\beta_{2}+d_{s-2}}\right)=\frac{a_{s-1}\beta_{1}}{c_{s-1}\beta_{1}+d_{s-1}},$
    ${\rm Tr}_{{K_{1}}/K_{0}}\left(\frac{a_{s-1}\beta_{1}}{c_{s-1}\beta_{1}+d_{s-1}}\right)=a_{s}c_{s}^{-1}.$
    It then follows from the transitivity of trace that
    ${\rm Tr}_{K_{s-1}}\left(\frac{\xi\beta_{s-1}-1}{\beta_{s-1}}\right)={\rm Tr}_{K_{0}}(a_{s}c_{s}^{-1})\neq 0.$
    Hence, (3.2) is still true for $i=s$, and then (3.2) holds for any $2\leq i\leq n$.
Therefore, each $h_{i}(X)$ is the minimal polynomial of $\beta_{i}$ over $K_{i-1}$. It gives that $[K_{i}:K_{i-1}]=p$ for all $1\leq i\leq n$, and then we have
$[K_{n}:K_{0}]=\prod_{i=1}^{n}[K_{i}:K_{i-1}]=p^{n}.$
From (3.2), one sees that $K_{n}=K_{0}(\beta_{n})$, so that $\beta_{n}$ is of degree $p^{n}$ over $K_{0}$. Note that $d_{n,g}(X)$ is a polynomial over $K_{0}$ of degree $p^{n}$ and $d_{n,g}(\beta_{n})=0$. Thus, we derive that $d_{n,g}(X)$ is irreducible over $K_{0}$, and the sufficiency is proved.
This completes the proof of Theorem 1.2. ∎
acknowledgment
This work was completed during the author’s academic visit to the Johann Radon Institute for Computational and Applied Mathematics (RICAM), Austrian Academy of Sciences. The author would like to express sincere gratitude to Professor Arne Winterhof at RICAM, for his careful reading of this paper and for providing valuable suggestions.